OpenStudy (anonymous):

5 + r = -19 someone pleaseee please help and show work please

OpenStudy (nubeer):

subtract 5 on both sides -5+5+r = -19-5 r = -19-5

OpenStudy (anonymous):

The point of these if to get the r on one side by itself. So we need to get rid of the 5 on the left. 5+r-5=-19-5 r=?

OpenStudy (anonymous):

what do i do? like im stuck with that whole question like where its numbers on both sides i forgot how to do that?

OpenStudy (anonymous):

You need to get all of the numbers on one side and all of the variables on another. The 5 + r is the part that is out of place because the -19 is fine. If you subtract 5 from 5+ r you get 5 + r -5 or 0 + r which is r. But because you subtract 5 from the left, you also have to from the right. So -19 - 5 is?

OpenStudy (anonymous):

14?

OpenStudy (nubeer):